OpenStudy (nfcfox):

lim as x approaches -3 where f(x) is (x^4 -3x^2 +19x+3)/(x+3)

OpenStudy (anonymous):

\[\frac{ x^{4}- 3x^{2} +19x+3 }{ x+3 } = \frac{ (x ^{3}-3x ^{2}+6x+1)(x+3) }{(x+3) } = (x ^{3}-3x ^{2}+6x+1)\]

OpenStudy (nfcfox):

Where are you getting the new numerator from.

OpenStudy (nfcfox):

Did you factor or something?

OpenStudy (anonymous):

\[(-3)^{3}-3(-3) ^{2}+6(-3)+1 \] is your answer..

OpenStudy (anonymous):

you have to divide.. to find out that form

OpenStudy (anonymous):

yeah i did factorize.. do you know the method?

OpenStudy (nfcfox):

So like long division? and can you explain end behavior to me by any chance?